TEXI2HTML(1) TEXI2HTML(1) NAME texi2html−aTexinfotoHTMLconverter SYNOPSIS texi2html[options]file texi2html-check[-verbose]files DESCRIPTION Texi2htmlconvertsthegivenTexinfofiletoasetofHTMLfiles.Ittriestohandlemostofthe Texinfocommands.Itcreateshypertextlinksforcross-references,footnotes... Italsotriestoaddlinksfromareferencetoitscorrespondingentryinthebibliography(ifany).It mayalsohandleaglossary(seethe−glossaryoption). Texi2htmlcreatesseveralfilesdependingonthecontentsoftheTexinfofileandonthechosen options(seeFILES). TheHTMLfilescreatedbytexi2htmlareclosertoTeXthantoInfo,that’swhytexi2htmlconverts @iftexsectionsandnot@ifinfoonesbydefault.Youcanreversethiswiththe−expandinfooption. OPTIONS −checkCheckthegivenfileandgivethelistofallthingsthatmaybeTexinfocommands. Thismaybeusedtochecktheoutputoftexi2htmltofindtheTexinfocommands thathavebeenleftintheHTMLfile. −expandinfo Expand@ifinfosections,not@iftexones. −glossaryUsethesectionnamed’Glossary’tobuildalistoftermsandputlinksintheHTML documentfromeachtermtowarditsdefinition. −invisiblename Usenametocreateinvisibledestinationanchorsforindexlinks.Thisisa workaroundforaknownbugofmanyWWWbrowsers,includingxmosaic. −IdirLookalsoindirtofindincludedfiles. −menuShowtheTexinfomenus;bydefaulttheyareignored. −monolithic Outputonlyonefile,includingthetableofcontentsandfootnotes. −numberNumberthesections. −splitchapter SplittheoutputintoseveralHTMLfiles(onepermainsection:chapter, appendix...). −splitnode SplittheoutputintoseveralHTMLfiles(onepernode). −usagePrintusageinstructions,listingthecurrentavailablecommand-lineoptions. −verboseGiveaverboseoutput.Canbeusedwiththe−checkoption. FILES Bydefaulttexi2htmlcreatesthefollowingfiles(foobeingthenameoftheTexinfofile): footoc.htmlThetableofcontents. foo.htmlThedocument’scontents. foofoot.htmlThefootnotes(ifany). Whenusedwiththe−splitoption,itcreatesseveralfiles(oneperchapterornode),named foon.html(nbeingtheindiceofthechapterornode),insteadofthesinglefoo.htmlfile. Whenusedwiththe−monolithicoption,itcreatesonlyonefile:foo.html 09/10/96 1 TEXI2HTML(1) TEXI2HTML(1) VARIABLES texi2htmlpredefinesthefollowingvariables:html,texi2html. ADDITIONALCOMMANDS texi2htmlimplementsthefollowingnon-Texinfocommands: @ifhtmlThisindicatesthestartofanHTMLsection,thissectionwillpassedthrough withoutanymodofication. @endifhtmlThisindcatestheendofanHTMLsection. VERSION Thisistexi2htmlversion1.51,09/10/96. Thelatestversionoftexi2htmlcanbefoundinWWW,cf.URL http://wwwcn.cern.ch/dci/texi2html/ AUTHOR ThemainauthorisLionelCons,CERNCN/DCI/UWS,Lionel.Cons@cern.ch.Manyotherpeople aroundthenetcontributedtothisprogram. COPYRIGHT ThisprogramistheintellectualpropertyoftheEuropeanLaboratoryforParticlePhysics(known asCERN).NoguaranteewhatsoeverisprovidedbyCERN.Noliabilitywhatsoeverisacceptedfor anylossordamageofanykindresultingfromanydefectorinaccuracyinthisinformationor code. CERN,1211Geneva23,Switzerland SEEALSO GNUTexinfoDocumentationFormat,HyperTextMarkupLanguage(HTML),WorldWideWeb (WWW). BUGS ThisprogramdoesnotunderstandallTexinfocommands(yet). TeXspecificcommands(normallyenclosedin@iftex)willbepassedunmodified. 09/10/96 2